OpenStudy (anonymous):

The sum of three consecutive integers is 108. Which equation would be used to solve this word problem? (HINT: The equation has been simplified some...) 3x + 3 = 108 3x = 108 3(x+3) = 108 3x + 1 = 108

OpenStudy (lgbasallote):

so which do you think?

OpenStudy (anonymous):

third one

OpenStudy (anonymous):

No..

OpenStudy (lgbasallote):

that's not quite right

OpenStudy (anonymous):

oops first one

OpenStudy (lgbasallote):

hint: let x = first number x+1 = second number x+ 3 = third number because they are consecutive

OpenStudy (anonymous):

Yes now you are right..

OpenStudy (lgbasallote):

that's x+2 = third number

OpenStudy (anonymous):

Thank You!

OpenStudy (lgbasallote):

the sum of those three is 108 so x + x + 1 + x + 2 = 108 3x + 3 = 108

OpenStudy (lgbasallote):

for the sake of solutions lol

OpenStudy (anonymous):

Thanks.. can you help me with one i just posted